# On vector-valued functional equations with multiple recursive terms

### v2

@@ 2 The Markov modulated $M/G/1$-type reflected autoregressive process
Customers arrive at time epochs $T_{1},T_{2},\ldots$, $T_{1}=0$, and service times are denoted by $S_{1},S_{2},\ldots$. If $X(t)=i$, arrivals occur according to a Poisson process with rate $\lambda_{i}>0$ and an arriving customer has a service time $S_{i}$, with cumulative distribution function (cdf) $B_{i}(.)$, Laplace-Stieltjes transform (LST) $\beta_{i}^{*}(s):=\int_{0}^{\infty}e^{-st}dB_{i}(t)$, $\bar{b}_{i}=-\beta_{i}^{*\prime}(0)$, $i=1,\ldots,N$, and $B^{*}(s):=diag(\beta_{1}^{*}(s),\ldots,\beta_{N}^{*}(s))$. We assume that given the state of the background process $\{X(t);t\geq 0\}$, $S_{1},S_{2},\ldots$ are independent and independent of the arrival process. Let $A_{n}=T_{n}-T_{n-1}$, $n=2,3,\ldots$, and $Y_{n}=X_{T_{n}}$, $n=1,2,\ldots$. We assume that such an arrival makes obsolete a fixed fraction $1-a_{i}$, given that $Y_{n}=i$, $i\in E$. Denote now by $R_{n}(X_{n})$ a random variable with support $\{a_{1},\ldots,a_{N}\}$, with $a_{i}\in(0,1)$, and such that $P(R_{n}(X_{n})=a_{i}|X_{n}=i)=1$.
Our focus is on the derivation of the transient distribution (in terms of Laplace-Stieltjes transform) of the process $\{(W_{n},T_{n});n=1,2,\ldots\}$ in which $\{T_{n};n=1,2,\ldots\}$ is an increasing time sequence generated by the input process, and $W_{n}$ denotes the workload in the system just before the $n$th customer arrival that takes place at $T_{n}$. This model generalizes the work in [20, Section 5], in which the author considered the case where the autoregressive parameter is independent of the state of the background state.
@@
$\begin{array}[]{l}E(e^{-sW_{n+1}-\eta T_{n+1}}1_{\{Y_{n+1}=j\}}|W_{1}=w)=E(e^{-s[R_{n}W_{n}+S_{n}-A_{n+1}]^{+}-\eta T_{n+1}}1_{\{Y_{n+1}=j\}})|W_{1}=w)\vskip 5.69054pt\\ =E((e^{-s[R_{n}W_{n}+S_{n}-A_{n+1}]}+1-e^{-s[R_{n}W_{n}+S_{n}-A_{n+1}]^{-}})e^{-\eta T_{n+1}}1_{\{Y_{n+1}=j\}})|W_{1}=w)\vskip 5.69054pt\\ =E(e^{-s[R_{n}W_{n}+S_{n}-A_{n+1}]-\eta T_{n+1}}1_{\{Y_{n+1}=j\}})|W_{0}=w)+E((1-e^{-s[R_{n}W_{n}+S_{n}-A_{n+1}]^{-}})e^{-\eta T_{n+1}}1_{\{Y_{n+1}=j\}}|W_{1}=w).\end{array}$
@@
$\begin{array}[]{r}E(e^{-s[R_{n}W_{n}+S_{n}-A_{n+1}]-\eta T_{n+1}}1_{\{Y_{n+1}=j\}}|W_{1}=w)=E(e^{-s[R_{n}W_{n}+S_{n}-A_{n+1}]-\eta(A_{n+1}+T_{n})}1_{\{Y_{n+1}=j\}}|W_{1}=w)\vskip 5.69054pt\\ =E(e^{-sR_{n}W_{n}-\eta T_{n}}e^{-sS_{n}-(\eta-s)A_{n+1}}1_{\{Y_{n+1}=j\}}|W_{1}=w)\vskip 5.69054pt\\ =\sum_{i=1}^{N}E(e^{-sa_{i}W_{n}-\eta T_{n}}1_{\{Y_{n}=i\}}|W_{1}=w)\beta_{i}^{*}(s)A_{i,j}(\eta-s)\end{array}$
@@
$\begin{array}[]{rl}E(e^{-sW_{n+1}-\eta T_{n+1}}1_{\{Y_{n+1}=j\}}|W_{1}=w)=&\sum_{i=1}^{N}E(e^{-sa_{i}W_{n}-\eta T_{n}}1_{\{Y_{n}=i\}}|W_{1}=w)\beta_{i}^{*}(s)A_{i,j}(\eta-s)\vskip 5.69054pt\\ &+E((1-e^{-s[R_{n}W_{n}+S_{n}-A_{n+1}]^{-}})e^{-\eta T_{n+1}}1_{\{Y_{n+1}=j\}}|W_{1}=w).\end{array}$ (10)
@@
since $V^{w}_{j}(r,0,\eta)=0$, $j=1,\ldots,N$. Thus, $C_{0,j}^{w}(r,\eta)=0$, so that $C_{0}^{w}(r,\eta)=(C_{0,1}^{w}(r,\eta),\ldots,C_{0,N}^{w}(r,\eta))^{T}=\mathbf{0}$.
@@
where $C^{w}_{i}(r,\eta):=(C^{w}_{i,1}(r,\eta),\ldots,C^{w}_{i,N}(r,\eta))^{T}$, $i=0,1,\ldots,N$, column vectors still have to be determined.
@@
$(-1)^{N}\prod_{i=1}^{N}\mu_{i}(\eta)[(I-r\Lambda(M^{T}(\eta))^{-1})\tilde{Z}^{w}(r,0,\eta)-r\widehat{p}]=\mathbf{0},$ (16)
since $\tilde{V}(r,0,\eta)=\mathbf{0}$, where $\mathbf{0}$, is $N\times 1$ column vector with all components equal to 0. Thus, $C^{w}_{0}(r,\eta)=\mathbf{0}$.
Note that (14) has the same form as eq. (50) in [2], although it is in matrix form. Moreover, (13) is the matrix analogue of [2, equation (48)], having, for $|r|<1$, $Re(\eta)\geq 0$, the form (see also [2, equation (2)])
$\mathbf{f}(r,s,\eta)=\mathbf{g}(r,s,\eta)\sum_{i=1}^{N}\tilde{P}^{(i)}\mathbf{f}(r,\alpha_{i}s,\eta)+\mathbf{K}(r,s,\eta),$ (17)
with $\mathbf{g}(r,s,\eta):=rA^{T}(\eta-s)B^{*}(s)$, $\mathbf{K}(r,s,\eta):=re^{-sw}\widehat{p}+\tilde{V}^{w}(r,s,\eta)$. Note that $\mathbf{g}(r,0,\eta)=r\Lambda(M^{T}(\eta))^{-1}=rA^{T}(\eta)$, $\mathbf{K}(r,0,\eta)=r\widehat{p}\neq\mathbf{0}$. Note that for $|r|<1$, $Re(\eta)\geq 0$ $|\mathbf{g}_{i,j}(r,0,\eta)|<1$, $|\mathbf{K}_{i,j}(r,0,\eta)|<1$, $i,j\in E$. Moreover, if for a matrix $M=(M_{i,j})_{i,j\in E}$, $M||=max_{1\leq i\leq N}\sum_{j=1}^{N}|M_{i,j}|$, then $\mathbf{g}(r,0,\eta)||<1$, $\mathbf{K}(r,0,\eta)||<1$.
It seems that a matrix generalization of [2, Theorem 2] can be applied to solve such kind of functional equations.
@@
$\begin{array}[]{rl}\tilde{Z}^{w}(r,s,\eta)=&\sum_{k=0}^{n}\sum_{i_{1}+\ldots+i_{N}=k}F_{i_{1},\ldots,i_{M}}(r,s,\eta)L^{w}(r,a_{i_{1},\ldots,i_{N}}(s),\eta)\vskip 5.69054pt\\ &+\sum_{i_{1}+\ldots+i_{N}=n+1}F_{i_{1},\ldots,i_{N}}(r,s,\eta)\tilde{Z}^{w}(r,a_{i_{1},\ldots,i_{N}}(s),\eta),\end{array}$ (19)
where $a_{i_{1},\ldots,i_{N}}(s)=a_{1}^{i_{1}}(a_{2}^{i_{2}}(\ldots(a_{N}^{i_{N}}(s))\ldots))$ and $a_{i}^{n}(s)$ denotes the $n$th iterate of $a_{i}(s)=a_{i}s$, with $a_{0,\ldots,0}(s)=s$, and the functions $F_{i_{1},\ldots,i_{N}}(r,s,\eta)$ are recursively defined by
$F_{i_{1},\ldots,i_{N}}(r,s,\eta)=r\sum_{k=1}^{N}F_{i_{1},\ldots,i_{k}-1,\ldots,i_{N}}(r,s,\eta)K^{w}(a_{i_{1},\ldots,i_{k}-1,\ldots,i_{N}}(s),\eta)\tilde{P}^{(k)},$
The next step is to investigate the convergence as $n\to\infty$. Note that $\tilde{P}^{(k)}||_{1}=1$, $k=1,\ldots,N$. Moreover,
[…]
$F_{i_{1},\ldots,i_{N}}(r,s,\eta)||\leq|r|^{n+1}\binom{i_{1}+\ldots+i_{N}}{i_{1},\ldots,i_{N}}C(\eta)(1+\epsilon)^{n+1}.$ (21)
@@
since the eigenvalues of $rA^{T}(\eta)$ are all strictly less than one when $Re(\eta)\geq 0$. So, as $n\to\infty$, by using (21), the second term of the right-hand side of (19) vanishes. Therefore, (19) becomes
$\begin{array}[]{rl}\tilde{Z}^{w}(r,s,\eta)=&\sum_{n=0}^{\infty}\sum_{i_{1}+\ldots+i_{N}=n}F_{i_{1},\ldots,i_{M}}(r,s,\eta)L^{w}(r,a_{i_{1},\ldots,i_{N}}(s),\eta).\end{array}$ (22)
@@
  • Substituting $s=\mu_{k}(\eta)$, $k=1,\ldots,N$ in (12) yield an $N^{2}\times N^{2}$ system of equations for the elements of $C^{w}_{j}(r,\eta)=(C^{w}_{j,1}(r,\eta),\ldots,C^{w}_{j,N}(r,\eta))^{T}$, $j=1,\ldots,N$:
@@
3 A Markovian fluid flow model with consumption
Consider a general stochastic fluid model with a single infinite capacity buffer where the buffer content $X(t)$ increases continuously. An external background process $J(t)$, $t\geq 0$ with a finite state space $E=\{1,2,\ldots,N\}$ affects the rate of the buffer increase. In particular, when $J(t)=i$, $i\in E$, the buffer content increases continuously with rate $r_{i}\in(-\infty,+\infty)-\{0\}$ (for convenience we assume that $r_{i}\neq 0$, $i\in E$). When $J(t)=i$ the process remains for an exponential amount of time with rate $q_{i}$, and then may move to any state $j$. At transition epochs, a fraction $1-a_{i}$ of the buffer content is instantaneously removed (i.e., is consumed). Let $T_{0}=0$ and $T_{1},T_{2},\ldots$ be the transition epochs of the process $\{J(t);t\geq 0\}$ with $T_{1}>0$. Let also $W_{0}=v$ and for $n=1,2,\ldots$, $W_{n}=X(T_{n})$ and $Y_{n}=J(T_{n}^{-})$ is the state of $\{J(t);t\geq 0\}$ just before transition epochs.
Let $A_{n}=T_{n}-T_{n-1}$ be the inter-jump time, $n=1,2,\ldots$, and define $R_{n}=\sum_{k=1}^{N}r_{k}1_{\{Z_{n}=k\}}$, $V_{n}=\sum_{k=1}^{N}a_{k}1_{\{Z_{n}=k\}}$. We then have the recursion:
@@ 3 A Markovian fluid flow model with consumption
Define for $Re(s)\geq 0$, $Re(\eta)\geq 0$, $|r|<1$,
$\begin{array}[]{rl}Z_{i,j}(r,s,\eta,v)=&\sum_{n=1}^{\infty}r^{n}E(e^{-sW_{n}-\eta T_{n}}1_{\{Y_{n}=j\}}|Y_{1}=i,W_{0}=v),\\ Z_{i}^{0}(s,\eta,v)=&E(e^{-sW_{1}-\eta T_{1}}1_{\{Y_{1}=i\}}|Y_{1}=i,W_{0}=v),\end{array}$
@@
$G_{i,j}(s,\eta)=E(e^{-(sr_{j}+\eta)A_{n+1}}1_{\{Y_{n+1}=j\}}|Y_{n}=i).$
Let $\mathbf{Z}(r,s,\eta,v)$, $\mathbf{G}(s,\eta)$, $\mathbf{V}(r,s,\eta,v)$ be $N\times N$ matrices with elements $Z_{i,j}(r,s,\eta,v)$, $G_{i,j}(s,\eta)$ and $V_{i,j}(r,s,\eta,v)$, respectively. Let also $\mathbf{Z}^{0}(s,\eta,v)=diag(Z_{1}^{0}(s,\eta,v),\ldots,Z_{N}^{0}(s,\eta,v))$. Note that
$\begin{array}[]{rl}Z_{i}^{0}(s,\eta,v)=&E(e^{-sW_{1}-\eta T_{1}}1_{\{Y_{1}=i\}}|Y_{1}=i,W_{0}=v)\\ =&\int_{0}^{\infty}e^{-s[v+r_{i}u]^{+}-\eta u}q_{i}e^{-q_{i}u}du.\end{array}$
@@
$\mathbf{G}(s,\eta)=P\mathbf{q}(\mathbf{q}+s\mathbf{r}+\eta I)^{-1},$
where $P$ is the one-step probability matrix of $\{Y_{n};n\in\mathbb{N}\}$, and $\mathbf{q}:=diag(q_{1},q_{2},\ldots,q_{N})$, $\mathbf{r}:=diag(r_{1},r_{2},\ldots,r_{N})$.
@@
$\begin{array}[]{l}E(e^{-sW_{n+1}-\eta T_{n+1}}1_{\{Y_{n+1}=j\}}|Y_{1}=i,V_{0}=v)=E(e^{-s[V_{n}W_{n}+R_{n+1}A_{n+1}]^{+}-\eta(T_{n}-A_{n+1})}1_{\{Y_{n+1}=j\}})|Y_{1}=i,V_{0}=v)\vskip 5.69054pt\\ =E(e^{-sV_{n}W_{n}-\eta T_{n}-(sR_{n+1}+\eta)A_{n+1}}1_{\{Y_{n+1}=j\}})|Y_{1}=i,V_{0}=v)\vskip 5.69054pt\\ +E((1-e^{-s[V_{n}W_{n}+R_{n+1}A_{n+1}]^{-}})e^{-\eta(T_{n}+A_{n+1})}1_{\{Y_{n+1}=j\}}|Y_{1}=i,V_{0}=v)\vskip 5.69054pt\\ =\sum_{k=1}^{N}E(e^{-sa_{k}W_{n}-\eta T_{n}}1_{\{Y_{n}=k\}})|Y_{1}=i,V_{0}=v)E(e^{-s(R_{n+1}+\eta)A_{n+!}}1_{\{Y_{n+1}=j\}})|Y_{n}=k)\vskip 5.69054pt\\ +E((1-e^{-s[V_{n}W_{n}+R_{n+1}A_{n+1}]^{-}})e^{-\eta(T_{n}+A_{n+1})}1_{\{Y_{n+1}=j\}}|Y_{1}=i,V_{0}=v)\vskip 5.69054pt\\ =\sum_{k=1}^{N}E(e^{-sa_{k}W_{n}-\eta T_{n}}1_{\{Y_{n}=k\}})|Y_{1}=i,V_{0}=v)G_{k,j}(s,\eta)\vskip 5.69054pt\\ +E((1-e^{-s[V_{n}W_{n}+R_{n+1}A_{n+1}]^{-}})e^{-\eta(T_{n}+A_{n+1})}1_{\{Y_{n+1}=j\}}|Y_{1}=i,V_{0}=v).\end{array}$
@@
$\mathbf{Z}(r,s,\eta,v)=r\sum_{k=1}^{N}\mathbf{Z}(r,a_{k}s,\eta,v)\tilde{P}^{(k)}\mathbf{G}(s,\eta)+\mathbf{L}(r,s,\eta,v),$ (29)
@@
$\mathbf{L}(r,s,\eta,v)=r\mathbf{Z}^{0}(s,\eta,v)+s\mathbf{C}^{(1)}(r,\eta,v)[s\mathbf{r}+\eta I+\mathbf{q}]^{-1},$
where $\mathbf{C}^{(1)}(r,\eta,v)$ an $N\times N$ matrix with $c_{i,j}^{(1)}(r,\eta,v)=0$, $i\in E$, $j\in R^{+}$, and $c_{i,j}^{(1)}(r,\eta,v)=0$, $i\in E$, $j\in R^{-}$ will be found by (28).

### v3

@@ 3.5 The Markov modulated $M/G/1$-type reflected autoregressive process
Customers arrive at time epochs $T_{1},T_{2},\ldots$, $T_{1}=0$, and service times are denoted by $S_{1},S_{2},\ldots$. If $X(t)=i$, arrivals occur according to a Poisson process with rate $\lambda_{i}>0$ and an arriving customer has a service time $S_{i}$, with cumulative distribution function (cdf) $B_{i}(.)$, LST $\beta_{i}^{*}(s):=\int_{0}^{\infty}e^{-st}dB_{i}(t)$, $\bar{b}_{i}=-\beta_{i}^{*\prime}(0)$, $i=1,\ldots,N$, and $B^{*}(s):=diag(\beta_{1}^{*}(s),\ldots,\beta_{N}^{*}(s))$. We assume that given the state of the background process $\{X(t);t\geq 0\}$, $S_{1},S_{2},\ldots$ are independent and independent of the arrival process. Let $A_{n}=T_{n}-T_{n-1}$, $n=2,3,\ldots$, and $Y_{n}=X_{T_{n}}$, $n=1,2,\ldots$. We assume that such an arrival makes obsolete a fixed fraction $1-a_{i}$, given that $Y_{n}=i$, $i\in E$. Denote now by $V_{n}(Y_{n})$ a random variable with support $\{a_{1},\ldots,a_{N}\}$, with $a_{i}\in(0,1)$, and such that $P(V_{n}(Y_{n})=a_{i}|Y_{n}=i)=1$.
Our focus is on the derivation of the transient distribution (in terms of LSTs) of the process $\{(W_{n},T_{n});n=1,2,\ldots\}$ in which $\{T_{n};n=1,2,\ldots\}$ is an increasing time sequence generated by the input process, and $W_{n}$ denotes the workload in the system just before the $n$th customer arrival that takes place at $T_{n}$. This model generalizes the work in [20, Section 5], in which the author considered the case where the autoregressive parameter is independent of the state of the background state.
@@
$\begin{array}[]{l}\mathbb{E}(e^{-sW_{n+1}-\eta T_{n+1}}1_{\{Y_{n+1}=j\}}|W_{1}=w)=\mathbb{E}(e^{-s[V_{n}W_{n}+S_{n}-A_{n+1}]^{+}-\eta T_{n+1}}1_{\{Y_{n+1}=j\}})|W_{1}=w)\vskip 5.69054pt\\ =\mathbb{E}((e^{-s[V_{n}W_{n}+S_{n}-A_{n+1}]}+1-e^{-s[V_{n}W_{n}+S_{n}-A_{n+1}]^{-}})e^{-\eta T_{n+1}}1_{\{Y_{n+1}=j\}})|W_{1}=w)\vskip 5.69054pt\\ =\mathbb{E}(e^{-s[V_{n}W_{n}+S_{n}-A_{n+1}]-\eta T_{n+1}}1_{\{Y_{n+1}=j\}})|W_{1}=w)+\mathbb{E}((1-e^{-s[V_{n}W_{n}+S_{n}-A_{n+1}]^{-}})e^{-\eta T_{n+1}}1_{\{Y_{n+1}=j\}}|W_{1}=w).\end{array}$
@@
$\begin{array}[]{r}\mathbb{E}(e^{-s[V_{n}W_{n}+S_{n}-A_{n+1}]-\eta T_{n+1}}1_{\{Y_{n+1}=j\}}|W_{1}=w)=\mathbb{E}(e^{-s[V_{n}W_{n}+S_{n}-A_{n+1}]-\eta(A_{n+1}+T_{n})}1_{\{Y_{n+1}=j\}}|W_{1}=w)\vskip 5.69054pt\\ =\mathbb{E}(e^{-sV_{n}W_{n}-\eta T_{n}}e^{-sS_{n}-(\eta-s)A_{n+1}}1_{\{Y_{n+1}=j\}}|W_{1}=w)\vskip 5.69054pt\\ =\sum_{i=1}^{N}\mathbb{E}(e^{-sa_{i}W_{n}-\eta T_{n}}1_{\{Y_{n}=i\}}|W_{1}=w)\beta_{i}^{*}(s)A_{i,j}(\eta-s)\end{array}$
@@
$\begin{array}[]{rl}\mathbb{E}(e^{-sW_{n+1}-\eta T_{n+1}}1_{\{Y_{n+1}=j\}}|W_{1}=w)=&\sum_{i=1}^{N}\mathbb{E}(e^{-sa_{i}W_{n}-\eta T_{n}}1_{\{Y_{n}=i\}}|W_{1}=w)\beta_{i}^{*}(s)A_{i,j}(\eta-s)\vskip 5.69054pt\\ &+\mathbb{E}((1-e^{-s[V_{n}W_{n}+S_{n}-A_{n+1}]^{-}})e^{-\eta T_{n+1}}1_{\{Y_{n+1}=j\}}|W_{1}=w).\end{array}$ (51)
@@
since $v^{w}_{j}(r,0,\eta)=0$, $j=1,\ldots,N$. Thus, $c_{0,j}^{w}(r,\eta)=0$, so that $C_{0}^{w}(r,\eta)=(c_{0,1}^{w}(r,\eta),\ldots,c_{0,N}^{w}(r,\eta))^{T}=\tilde{0}$.
@@
where $C^{w}_{i}(r,\eta):=(c^{w}_{i,1}(r,\eta),\ldots,c^{w}_{i,N}(r,\eta))^{T}$, $i=0,1,\ldots,N$, the column vectors still have to be determined.
@@
$(-1)^{N}\prod_{i=1}^{N}\mu_{i}(\eta)[(I-r\Lambda(M^{T}(\eta))^{-1})\tilde{Z}^{w}(r,0,\eta)-r\widehat{p}]=\tilde{0}$, (57)
since $\tilde{V}(r,0,\eta)=\tilde{0}$, where $\tilde{0}$, is an $N\times 1$ column vector with all components equal to 0. Thus, $C^{w}_{0}(r,\eta)=\tilde{0}$.
@@
$\begin{array}[]{rl}\tilde{Z}^{w}(r,s,\eta)=&\sum_{k=0}^{n}\sum_{i_{1}+\ldots+i_{N}=k}F_{i_{1},\ldots,i_{M}}(r,s,\eta)L^{w}(r,\alpha_{i_{1},\ldots,i_{N}}(s),\eta)\vskip 5.69054pt\\ &+\sum_{i_{1}+\ldots+i_{N}=n+1}F_{i_{1},\ldots,i_{N}}(r,s,\eta)\tilde{Z}^{w}(r,\alpha_{i_{1},\ldots,i_{N}}(s),\eta),\end{array}$ (59)
where $\alpha_{i_{1},\ldots,i_{N}}(s)=\alpha_{1}^{i_{1}}(\alpha_{2}^{i_{2}}(\ldots(\alpha_{N}^{i_{N}}(s))\ldots))$ and $\alpha_{i}^{n}(s)$ denotes the $n$th iterate of $\alpha_{i}(s)=a_{i}s$, with $\alpha_{0,\ldots,0}(s)=s$, and the functions $F_{i_{1},\ldots,i_{N}}(r,s,\eta)$ are recursively defined by
$F_{i_{1},\ldots,i_{N}}(r,s,\eta)=r\sum_{k=1}^{N}F_{i_{1},\ldots,i_{k}-1,\ldots,i_{N}}(r,s,\eta)K^{w}(\alpha_{i_{1},\ldots,i_{k}-1,\ldots,i_{N}}(s),\eta)\tilde{P}^{(k)},$
with $F_{0,\ldots,0}(r,s,\eta):=I$, and $F_{i_{1},\ldots,i_{N}}(r,s,\eta)=O$ (that is, the zero matrix) if one of the indices is equal to 1-1. It is clear that (58) is of the form (2), with $U(s,\eta):=K^{w}(s,\eta)$, and $K(r,s,\eta):=L^{w}(r,s,\eta)$, and where the contractions $\alpha_{i}(s)$ have a common fixed-point $a=0$. Moreover, as $i_{1}+\ldots+i_{N}\to\infty$, $K^{w}(\alpha_{i_{1},\ldots,i_{N}}(s),\eta)\to A^{T}(\eta)$, and $L^{w}(r,\alpha_{i_{1},\ldots,i_{N}}(s),\eta)\to r\widehat{p}$.
@@
since the eigenvalues of $rA^{T}(\eta)$ are strictly less than one when $Re(\eta)\geq 0$. So, by applying Theorem 2 we have
$\begin{array}[]{rl}\tilde{Z}^{w}(r,s,\eta)=&\sum_{n=0}^{\infty}\sum_{i_{1}+\ldots+i_{N}=n}F_{i_{1},\ldots,i_{M}}(r,s,\eta)L^{w}(r,\alpha_{i_{1},\ldots,i_{N}}(s),\eta).\end{array}$ (60)
@@
  • Substituting $s=\mu_{k}(\eta)$, $k=1,\ldots,N$ in (53) yields a $N^{2}\times N^{2}$ system of equations for the elements of $C^{w}_{j}(r,\eta)=(c^{w}_{j,1}(r,\eta),\ldots,c^{w}_{j,N}(r,\eta))^{T}$, $j=1,\ldots,N$:
@@
3.6 A Markovian fluid flow model with consumption
Consider a general stochastic fluid model with a single infinite capacity buffer where the buffer content $X(t)$ increases continuously. An external background process $J(t)$, $t\geq 0$ with a finite state space $E=\{1,2,\ldots,N\}$ affects the rate of the buffer increase. In particular, when $J(t)=i$, $i\in E$, the buffer content continuously increases with rate $r_{i}\in(-\infty,+\infty)-\{0\}$ (for convenience, we assume that $r_{i}\neq 0$, $i\in E$). When $J(t)=i$ the process remains for an exponential amount of time with rate $q_{i}$, and then may move to any state $j$. At transition epochs, a fraction $1-a_{i}$ of the buffer content is instantaneously removed (i.e., it is consumed). Let $T_{0}=0$ and $T_{1},T_{2},\ldots$ be the transition epochs of the process $\{J(t);t\geq 0\}$ with $T_{1}>0$. Let also $W_{0}=v$ and for $n=1,2,\ldots$, $W_{n}=X(T_{n})$ and $Y_{n}=J(T_{n}^{-})$ be the state of $\{J(t);t\geq 0\}$ just before transition epochs.
Let $A_{n}=T_{n}-T_{n-1}$ be the inter-jump time, $n=1,2,\ldots$, and define $R_{n}=\sum_{k=1}^{N}r_{k}1_{\{Y_{n}=k\}}$, $V_{n}=\sum_{k=1}^{N}a_{k}1_{\{Y_{n}=k\}}$. We then have the following recursion:
@@ 3.6 A Markovian fluid flow model with consumption
Define for $Re(s)\geq 0$, $Re(\eta)\geq 0$, $|r|<1$, $i,j\in E$
$\begin{array}[]{rl}z_{i,j}(r,s,\eta,v)=&\sum_{n=1}^{\infty}r^{n}\mathbb{E}(e^{-sW_{n}-\eta T_{n}}1_{\{Y_{n}=j\}}|Y_{1}=i,W_{0}=v),\vskip 5.69054pt\\ z_{i}^{0}(s,\eta,v)=&\mathbb{E}(e^{-sW_{1}-\eta T_{1}}1_{\{Y_{1}=i\}}|Y_{1}=i,W_{0}=v),\end{array}$
@@
$g_{i,j}(s,\eta)=\mathbb{E}(e^{-(sr_{j}+\eta)A_{n+1}}1_{\{Y_{n+1}=j\}}|Y_{n}=i).$
Let $Z(r,s,\eta,v)$, $G(s,\eta)$, $V(r,s,\eta,v)$ be $N\times N$ matrices with elements $z_{i,j}(r,s,\eta,v)$, $g_{i,j}(s,\eta)$ and $v_{i,j}(r,s,\eta,v)$, respectively. Let also $Z^{0}(s,\eta,v)=diag(z_{1}^{0}(s,\eta,v),\ldots,z_{N}^{0}(s,\eta,v))$. Note that
$\begin{array}[]{rl}z_{i}^{0}(s,\eta,v)=&E(e^{-sW_{1}-\eta T_{1}}1_{\{Y_{1}=i\}}|Y_{1}=i,W_{0}=v)=\int_{0}^{\infty}e^{-s[v+r_{i}u]^{+}-\eta u}q_{i}e^{-q_{i}u}du.\end{array}$
@@
$G(s,\eta)=P\widehat{Q}(\widehat{Q}+\widehat{R}+\eta I)^{-1},$
where $P$ is the one-step probability matrix of $\{Y_{n};n\in\mathbb{N}\}$, and $\widehat{Q}:=diag(q_{1},q_{2},\ldots,q_{N})$, $\widehat{R}:=diag(r_{1},r_{2},\ldots,r_{N})$.
@@
$\begin{array}[]{l}\mathbb{E}(e^{-sW_{n+1}-\eta T_{n+1}}1_{\{Y_{n+1}=j\}}|Y_{1}=i,W_{0}=v)=\mathbb{E}(e^{-s[V_{n}W_{n}+R_{n+1}A_{n+1}]^{+}-\eta(T_{n}-A_{n+1})}1_{\{Y_{n+1}=j\}})|Y_{1}=i,W_{0}=v)\vskip 5.69054pt\\ =\mathbb{E}(e^{-sV_{n}W_{n}-\eta T_{n}-(sR_{n+1}+\eta)A_{n+1}}1_{\{Y_{n+1}=j\}})|Y_{1}=i,W_{0}=v)\vskip 5.69054pt\\ +\mathbb{E}((1-e^{-s[V_{n}W_{n}+R_{n+1}A_{n+1}]^{-}})e^{-\eta(T_{n}+A_{n+1})}1_{\{Y_{n+1}=j\}}|Y_{1}=i,W_{0}=v)\vskip 5.69054pt\\ =\sum_{k=1}^{N}\mathbb{E}(e^{-sa_{k}W_{n}-\eta T_{n}}1_{\{Y_{n}=k\}})|Y_{1}=i,W_{0}=v)\mathbb{E}(e^{-s(R_{n+1}+\eta)A_{n+1}}1_{\{Y_{n+1}=j\}})|Y_{n}=k)\vskip 5.69054pt\\ +\mathbb{E}((1-e^{-s[V_{n}W_{n}+R_{n+1}A_{n+1}]^{-}})e^{-\eta(T_{n}+A_{n+1})}1_{\{Y_{n+1}=j\}}|Y_{1}=i,W_{0}=v)\vskip 5.69054pt\\ =\sum_{k=1}^{N}\mathbb{E}(e^{-sa_{k}W_{n}-\eta T_{n}}1_{\{Y_{n}=k\}})|Y_{1}=i,W_{0}=v)g_{k,j}(s,\eta)\vskip 5.69054pt\\ +\mathbb{E}((1-e^{-s[V_{n}W_{n}+R_{n+1}A_{n+1}]^{-}})e^{-\eta(T_{n}+A_{n+1})}1_{\{Y_{n+1}=j\}}|Y_{1}=i,W_{0}=v).\end{array}$
@@
$Z(r,s,\eta,v)=r\sum_{k=1}^{N}Z(r,a_{k}s,\eta,v)\tilde{P}^{(k)}G(s,\eta)+L(r,s,\eta,v),$ (67)
@@
$L(r,s,\eta,v)=rZ^{0}(s,\eta,v)+sC^{(1)}(r,\eta,v)[s\widehat{R}+\eta I+\widehat{Q}]^{-1},$
where $C^{(1)}(r,\eta,v)$ an $N\times N$ matrix with $c_{i,j}^{(1)}(r,\eta,v)=0$, $i\in E$, $j\in R^{+}$, and $c_{i,j}^{(1)}(r,\eta,v)$, $i\in E$, $j\in R^{-}$ are obtained by (66).